OpenStudy (anonymous):

Determine if (-2/3,5) is a solution to -3x+2y=-4 . Please show all of your work

OpenStudy (anonymous):

Plug the x and y values into the equation and determine if it's true. -3(-2/3)+2(5)=-4

OpenStudy (anonymous):

Did you work it out or do you need my help?

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

Alright solve the equation i made earlier -3(-2/3)+2(5)=-4 2+10=-4 12=-4 <------ not true so therefore (-2/3, 5) is not a solution.

OpenStudy (anonymous):

thank you so much I understand it now

OpenStudy (anonymous):

yw